28. (a) Given the equation

x^2+qx+2a^2-b^2+ab=0

i) Factorize the independent term of the equation /3mrks

ii) After determining the product of roots of (E), deduce the roots given that one of them is

a +b /3mrks


iii) Deduce q in terms of a and b. /2mrks

(b) Determine two numbers whose S = 4a -b and P 3a^2- 2b^2- 5ab /4mrks


(c) Given the equation 3x^2+ kx + 8 = 0. Determine the roots of this equation if their ratio is 1.5.


Deduce the value of k.


1
Expert's answer
2021-09-06T16:42:59-0400
"x^2+qx+2a^2-b^2+ab=0"

i)


"2a^2-b^2+ab=a^2-b^2+a^2+ab"

"(a-b)(a+b)+a(a+b)=(2a-b)(a+b)"

ii)


"x_1x_2=\dfrac{(2a-b)(a+b)}{1}=(2a-b)(a+b)"

If "x_1=a+b," then "x_2=2a-b."

"\{ a+b, 2a-b\}"


iii)


"x_1+x_2=-\dfrac{q}{1}=-q"

"q=-(a+b+2a-b)"

"q=-3a"

(b)


"x_1+x_2=4a-b"

"x_1x_2=3a^2-2b^2-5ab"

"x^2-(4a-b)x+3a^2-2b^2-5ab=0"

"D=(4a-b)^2-4(3a^2-2b^2-5ab)"

"=16a^2-8ab+b^2-12a^2+8b^2+20ab"

"=4a^2+12ab+9b^2=(2a+3b)^2"

"x=\dfrac{-(-(4a-b))\pm\sqrt{(2a+3b)^2}}{2}"

"=\dfrac{4a-b\pm(2a+3b)}{2}"

"x_1=\dfrac{4a-b-(2a+3b)}{2}=\dfrac{4a-b-2a-3b}{2}=a-2b"

"x_2=\dfrac{4a-b+(2a+3b)}{2}=\dfrac{4a-b+2a+3b}{2}=3a+b"

(c)


"3x^2+ kx + 8 = 0"

"x_1=1.5x_2"

"x_1x_2=\dfrac{8}{3}"

"1.5x_2^2=\dfrac{8}{3}"

"x_2^2=\dfrac{16}{9}"

"x_2=-\dfrac{4}{3}, x_1=-2, k=x_1+x_2=-\dfrac{10}{3}"

Or


"x_2=\dfrac{4}{3}, x_1=2, k=x_1+x_2=\dfrac{10}{3}"